[[Continuum/NIKA2/Main|Back to the NIKA2 main page]] <<BR>>

= Daily Reports =

Observers on-site: Maite Beltrán, Anaëlle Maury (until Friday), Andrés Zuleta, Rodrigo Álvarez and Angel (AoD). Dave, Carlos and Jean supporting on-site (Tuesday, maintenance slot).

Operators: Gloria (day shift) and Sergio (night shift)

'''NOTE by SB''': if possible, please take some total power beam maps (Uranus at best); and some pol beam maps on Uranus too. 

 
== Jan 28th, Tuesday ==

14:00-16:00; Angel. No system check. NIKA2 and subsystems in good shape. Very bad weather (collateral effects of the storm Herminia; Ivo is coming). EMIR checks of high-frequency signal of external calibrator pending. 

16:00-19:00; Rodrigo. Clear skies and tau_225GHz~0.15 after the last storm. Taumeter's external box wiped. Deicing system at work.

19:00-02:00: Anaëlle & Andrés. Sergio inspected the telescope from the platform (~19:30 LT). Ambient temp. -10C. Hard ice block on the upper third of the subreflector. A thumb-thick ice layer on the lower-half of M1. Telescope stopped. 

== Jan 29th, Wednesday ==

02:00-06:00; Angel: Ice on M1 and M2. Deicing active. Telescope stopped.

06:00-13:00; Maite: Antenna still stopped at 06:00. Deicing active all night but not properly working in some sectors (e.g., 8 and 3). Sergio moved the antenna to sun rising direction and waiting. Opacity at 0.3 and rising. \\

 * 08:00: Cloudy skies, opacity 0.3. Waiting. Still ice on the bottom side of the antenna. 

 * 12:45: Ice almost melted. Wind gusts 15 m/s. Opacity 0.26.  
 
 * 13:00: We will start with test on EMIR.

13:00-19:00: Rodrigo: tests using t07, pointing, onoff_psw. Antenna parked due to high winds (14:00).

 * 16:00. Testing pointing t07, bad results, parking antenna.

19:00-02:00: Anaëlle & Andrés. Snow and wind. Telescope parked.

== Jan 30th, Thursday ==

02:00-06:00; Angel: Snow storm, high wind with a brief stop at ~01:30 (LT). Ambient temp. -9C. Deicing fully working, except for sector 8 (see below) mainly, and some interleaved sections of the sub-reflector cell (as revealed by thermal camera shots, MC). Waiting for a sunny morning for deicing. Light fog at the end of the shift.  

[[attachment:image_2025_01_30T04_18_27_914Z.png|{{attachment:image_2025_01_30T04_18_27_914Z.png|ExCal|height=300}}|target="_blank"]]

06:00-13:00; Maite: Stopped snowing but we are in the clouds. Deicing working, waiting for the Sun. Temperature -10C.

 * 12:00: Still plenty of ice on antenna surface. The wind does not allow to use the platform to clean the secondary.   

13:00-19:00; Rodrigo. Antenna still parked.

19:00-02:00; Anaëlle & Andrés: Fog and wind. Telescope full of ice, impossible to clean: parked.

== Jan 31th, Friday ==

02:00-06:00; Angel. Wind speed dropped enough to allow the cleaning of secondary cell, sub-reflector and legs from the platform. Thanks, Sergio. Telescope cleared, finally (02:45). Deicing switched off. Wind speed 12-13 m/s (from OSN anemometer; ours is frozen, as well as the vane). Relative humidity sensor is frozen too (reading blocked at 100%). Tau_225GHz~0.16. 

 * Tuning E0+E2 with EMIR for testing beam geometry after deicing and the external calibrator. FWHM 12" x 13" (1.2 mm) and 26" x 27" (3.3 mm). Some unexplained spikes on spectra of the EC tests (see the Operator report). DJ suggests a new test but with single Rx configuration, among other.

 * 05:00: Switching to NIKA2, total power. Bad KIDs: 24/3/35. Tau_225GHz~0.17. Quite stable: p2p variations +/- 1 Jy/b (2 mm) along focus sequences.  Project 189-24, 1 repetition.
 
06:00-13:00; Maite. Clear skies, opacity ~0.2. Finishing 189-24 in TP. Starting polarization @ 07:15

 * 07:30: Skydip (scan 20250131s55), pointing and beam map-pol on 3C286 (s57). Beam 17.6"x17.7". No focus needed.  Calib_1scan on 3C286.

 * 08:50: Starting with project 145-24. Source G16.59. Opacity 0.15. 1 Repetition.

 * 10:15: Pointing+calib_1 Focus+Pointing+Calib_1 on 3C286, elev. 33

 * 10:45: Project 145-24, another repetition on G16.59. Opacity ~0.15.

 * 12:15: Pointing, focus, and calib_1 on 3C454. Opacity 0.12.

 * 12:40: Problems with the spindles of antenna (first since we started observing). Stopped observations for a while. Antenna stowed for security.

13:00-19:00; Rodrigo: Pointing on Uranus, and 3C454.

 * 17:00: observing project 135-54, last iteration of NGC7538 done. The anemometer is working again.  

 * 18:00: pointing and calib1_scan 3C454, calibrations on CRL2688, MWC349, NGC7027 (scans 140-142). Skydip (scan s143). Pointing on Uranus, then moving to 128-24

19:00-02:00; Andres: Hello world. Clear skies, stable atmosphere. Opacity ~0.11. Starting with project 128-24, source L1551.

 * 21:50: Last iteration of L1551 done. Changing to nikaw-24 account.

 * 22:10: Calibrations finished. Moving to project 015-17, source NGC1333, 2 reps.

 * 00:30: Changing to nikaw-24 account for calibrations.

 * 00:40: Starting with project 136-24.
 
== Feb 1st, Saturday ==

02:00-06:00; Angel. Tau_225GHz~0.18, quite stable atmosphere. 

 * Finishing 3rd round on L1482 (136-24), when target reached El. = 30 deg. + end calibration. 

 * Next science target (project 189-24 only for current LST) is too low for observing. 

 * Taking this opportunity to test the external calibrator (EC), this time with single Rx configuration. Three single scans with EC = off-on-off, for harmonics 15 (232.5 GHz: scans 17, 19, 21) and 17 (263.5 GHz: scans 25, 27, 29). The effect of the EC is clear:

[[attachment:s17.png|{{attachment:s17.png|ExCal|height=200}}|target="_blank"]]
[[attachment:s19.png|{{attachment:s19.png|ExCal|height=200}}|target="_blank"]]
[[attachment:s21.png|{{attachment:s21.png|ExCal|height=200}}|target="_blank"]] <<BR>>


[[attachment:s25.png|{{attachment:s25.png|ExCal|height=200}}|target="_blank"]]
[[attachment:s27.png|{{attachment:s27.png|ExCal|height=200}}|target="_blank"]]
[[attachment:s29.png|{{attachment:s29.png|ExCal|height=200}}|target="_blank"]] <<BR>>

 * 03:30: Switching to NIKA2 again, polarimetry. Bad KIDs: 21/3/40. Tau_225GHz~0.17. Quite stable: p2p variations +/- 1 to 2 Jy/b (2 mm) along focus sequences. Beam map-pol on 3C286 (scan 20250201s38).

 * 04:30: Changing to NIKA2, total power. ad KIDs: 25/5/39. Tau_225GHz~0.18. Very stable. Project 189-24, 1 repetition on CTRS27.

06:00-13:00; Maite. Continuing with project 189-24. Opacity ~0.2. Clear skies. Outside temperature -5C. 

 * 07:15: Project 189-24: Scan 1/4 of OTF map on CRTS27 (20250201s60) shows some spikes in the signal plots. Checked and one KID having a weird behavior. It doesn't seem to affect the final map.  Scan 4/4 (20250201s63) has some central spikes.

 * 08:15: Starting NIKA2 in Pol. mode. Band KIDs number, 18, 3, and 39. Quite stable. Tau < 0.2. Pointing, Focus and calib_1 on 3C286.  
 
 * 08:45: Starting project 145-24, source G16.59. Tau ~ 0.2. Some clouds in the sky. 1 repetition.

 * 09:10: Pointing (20250201s82) and calib_1scan (20250201s83) on 3C286. No focus needed. Tau slightly rising (~0.25) 

 * 09:30: Continuing with project 145-24, source G16.59. 1 repetition. Tau a bit high (0.20-0.25)  

 * 10:50: Pointing (20250201s93) and calib_1scan (20250201s93) on 3C454. Then focus (20250201s94-98), pointing (20250201s99), and calib_1scan (20250201s100) again. 

 * 11:30: Starting with project 128-24, source B335. 1 repetition. Source at elev. ~ 60 degrees. Clear skies, Tau ~ 0.2.  

13:00-19:00: Rodrigo. only first round of B335 given we made calibrations in between. After calibration we continued with the second round of B335

 * 14:30: project 135-24: pointing and then one repetition on IRAS21078 

 * 15:30: 3C454 pointing and calib1_scan, pointing on 2037+511.

 * 15:50: One repetition of IRAS21078.

 * 16:45: 3C454 pointing, calib_1scan, then pointing in 2037+511

 * 17:00: last repetition of IRAS21078. Source finished

 * 17:55: Pointing and calibration on MWC349, then calibration on CRL2688

 * 18:30: Starting project 166-24. Calibration and pointing on 3C454. Then pointing on NGC7538

 * 18:48: Start observations of CasA, two repetitions of observe_CasA script.

19:00-02:00: Andres. Continuing with project 166-24 

 * 20:05: Changing to nikaw-24 account. Pointing in Uranus 

 * 20:15: Back to project 166-24. calib_1scan in Uranus, pointing in 2254+617, then moving to CasA for observations (two repetitions of the observe_CasA script). Finished with pointing and calib_1scan in Uranus.

 * 21:50: Switching to nikaw-24. Focus correction in Uranus, pointing, and calib_1scan.

 * 22:20: Moving to project 015-17. calib_1scan. Uranus is close to the source, we moved directly to NGC1333, 1 rep. before a new calib_1scan on Uranus.

 * 23:20: Switching to nikaw-24. Pointing/focus on Uranus. Beam map-pol on Uranus (20250201s210). Moved to 3C84 for pointing, focus, and pointing.

 * 23:30: Moving to project 136-24: calib_1scan in 3C84. Pointing in CRL618. Observing L1482. Spindle error in scan 222. Fixed quickly (SA).

== Feb 2nd, Sunday ==

02:00-06:00; Angel. Tau_225GHz < 0.1, quite good stability. Ambient temp. -6C. 

 * Closing observations of L1482 (project 136-24) with a calib_1scan on Mars. 

 * Changing to nikaw-24. A calib_1scan (20240202s15) on alpha Ori. Beam map-pol on 3C286 (20250202s24). Skydip (20250202s25).

 * 03:50: switching to total power. DAQ restarted, bad KIDs: 13,3,40. 

 * Moving to project 189-24, target CTRS27. Tau_225 GHz = 0.09. Quite stable atmosphere: p2p vars. < 1 Jy/beam (2 mm).

 * 04:02 (LT) Spindle error, scan 29, cancelled. Fixed quickly right after recording main ACU parameters with Scope (SA).

 * One rep. on CTRS27 (189-24). A second one was started.

06:00-13:00; Maite. Clear skies. Extremely good opacity (< 0.1). Outside temperature -7C. Finishing repetition on CTRS27 (189-24). Spindle error at end (~ subscan 40) of scan 45 (map direction 90). Finishing the scan and resetting. Last map done (direction 135). 
   
 * 07:05: Starting a new repetition on CRTS27 (project 189-24). Tau < 0.1. Spindle problem at scan 50 in subscan ~ 11 (map direction 90). Last map (direction 135) not done. Start with direction 135 on the next session.
 
 * 08:10: Changing to nikaw-24. Switching to polarization. Clear skies, tau ~ 0.1. Bad KIDs 11, 3, 38. Problems starting the polarization mode. Restarting the server. Bad KIDs 8, 3, and 39.

 * 08:40: Pointing (20250202s52), focus (20250202s53-57), focus (20250202s58-62), pointing (20250202s63), calib_1scan (20250202s64) on 3C286. Tau < 0.1  

 * 09:10: Starting project 145-24, source G16.59. Tau ~0.12. 1 repetition.   

 * 10:40:  Pointing (20250202s72), calib_1scan (20250202s73) on 3C286. Tau ~0.1. Changing to nika2-w and doing focus (s74-78), but problems with the creation of imbfits. Operator stops the acquisition. Bad KIDs 10, 3, 39

 * 11:12: Again focus (20250202s79-83) and calib_1scan (20250202s84) on 3C286 (a bit low in elevation, 25). Tau < 0.09

 * 11:30: Project 145-24, source G16.59. 1 repetition.  Tau <0.1.

 * 12:50: Pointing (20250202s92) and calib_1scan (20250202s93) on 3C454. Elevation 52. Tau < 0.1 

13:00-19:00: Rodrigo. starting with a focus and pointing on 3C454.

 * 13:45: Observation of B335, one round first.      

 * 13:55: Sudden arrival of cloud banks. Tiny ice crystals can be seen backlight. The halo around the sun demonstrates it. Disappeared in ~15 min. Atmosphere gets a bit unstable. 
         Last scans (108-110) of the first round are likely affected by atmosphere instabilities.

[[attachment:sunhalo_20250202.jpeg|{{attachment:sunhalo_20250202.jpeg|ExCal|height=300}}|target="_blank"]]

 * 14:20: pointing and calib_1scan on 3C454.

 * 14:37: Pointing on 1923+210.

 * 14:42: one round of B335

 * 15:11: pointing 3C454 and calib_1scan

 * 15:23: pointing on 1923+210

 * 15:30: one round of B335

 * 16:00: pointing and calib_1scan 3C454

 * 16:20: '''Switched to total power''', DAQ restarted, bad KIDs: KIDS 14, 3,38

 * 16:30: pointing and focus on 3C454.

 * 16:45: one repetition of observe_A2390, scan 2 has missing most KIDs in one subunit of array 2, after completing scan 3 we send the scan with angle 120, completing the observe_A2390_half instead

 * 17:50: Switch to nikaw-24 account for calibrations on 3C454, MWC349 (calib_1scan, 20250202s156), CRL2688 (calib_1scan, s157), NGC7027 (calib_1scan, s158)

 * 18:10: Moving to project 114-24. Starting one repetition on Cep_1.

19:00-02:00: Andres. Continued with Cep_1. Done successfully.

 * 19:45: Switched to nikaw-24 account. Performed pointing, focus, pointing, beam map (s173), skydip (s174), pointing, and calib_1scan in Uranus. '''All in total power.'''

 * 20:10: '''Switched to polarimetry''': pointing, focus, pointing, beam map-pol (s184), skydip (s185), and pointing. Last pointing was not created and we stopped for a bit (22:12 - 22:32). Pointing showed quite unstable conditions. We ran another pointing. Look at the at the opacity and array 2 timelines attached below.

[[attachment:Screenshot from 2025-02-02 22-45-26.png|{{attachment:Screenshot from 2025-02-02 22-45-26.png|ExCal|height=300}}|target="_blank"]]
[[attachment:Screenshot from 2025-02-02 22-46-04.png|{{attachment:Screenshot from 2025-02-02 22-46-04.png|ExCal|height=300}}|target="_blank"]] <<BR>>

 * 23:00: Moving to EMIR, project 186-24. After 3 pointing tries (under project ''test''), no reasonable values were found. Very unstable weather conditions.

 * 00:25: It started to snow. Parking the telescope.

== Feb 3rd, Monday ==

02:00-06:00; Angel. Still snowing; tau_225GHz > 0.6 with large amplitudes. No wind. Telescope stopped.

06:00-11:00: Project 087-23 (POLAMI) scheduled. Snowfall continues. Deicing ON.

End of the 3rd W24 NIKA2 pool week. Holography campaign starts.